$SPAD/src/input bug10312.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool bug10312.output
)set message test on
)set message auto off
)clear all
--S 1 of 2
p:=(1/2+n)::UTS(FRAC INT, 'n, 0)
--R
--R
--R        1
--R   (1)  - + n
--R        2
--R                          Type: UnivariateTaylorSeries(Fraction(Integer),n,0)
--E 1
--S 2 of 2
(p**(-1))$UTS(FRAC INT, 'n, 0)
--R
--I   Compiling function G1473 with type Integer -> Boolean
--R
--R   (2)
--R                2      3      4      5       6       7       8        9
--R     2 - 4n + 8n  - 16n  + 32n  - 64n  + 128n  - 256n  + 512n  - 1024n
--R   +
--R          10      11
--R     2048n   + O(n  )
--R                          Type: UnivariateTaylorSeries(Fraction(Integer),n,0)
--E 2
)spool
)lisp (bye)

———-

References^{[1] nothing}

4